OpenStudy (anonymous):

How do I find an anti-derivative that is the inverse of the sum rule? (f+g)'(x) = f'(x) + g'(x)?

OpenStudy (anonymous):

dont understand... there is sum rule for integrals too. its the same

OpenStudy (anonymous):

sorry, anti-derivative rule...

OpenStudy (anonymous):

\[\int (a(x)+b(x))dx=\int a(x) dx+\int b(x) dx\]

OpenStudy (anonymous):

ok, cool